# Seeing and Seeing Through the Glass: Real and Synthetic Data for Multi-Layer Depth Estimation

### v1

@@ 1 Introduction
Our real-world benchmark consists of 1,500 images of transparent objects collected from diverse environments, including households, retail spaces, restaurants, laboratories, urban environments, and art installations, under various lighting conditions. Since ground-truth numerical depth cannot be accurately obtained for transparent objects, let alone multi-layer depth, we turn to relative depth annotations instead. Relative depth provides rich information about 3D structures and serves as an effective evaluation metric. Moreover, human annotators excel at determining relative depth, as they can reliably judge which of two points is closer to the camera and provide accurate annotations. In total, we generate 10.2M tuples for relative depth annotations. A gallery of our benchmark along with sample annotations is shown in Fig. 1. All images and validation annotations are available under CC0. Our benchmark is highly challenging for state-of-the-art depth estimation methods [48, 49, 4, 15, 19, 3, 26, 42, 17, 31], even when evaluated on the simplified task of predicting only the first visible layer. For example, Metric3D V2 [17] achieves just 55.14% quadruplet accuracy, while Depth Anything V2 [49], the most accurate among them, reaches only 70.43%.
@@
  • We propose a real-world multi-layer depth benchmark LayeredDepth for transparent objects, consisting of 1,500 CC0 images of diverse scenes and 10.2M relative depth tuples. Our evaluation of state-of-the-art depth estimation methods on our benchmark reveals that they struggle significantly with transparent objects.
@@ 4.2 Annotation Acquisition
The annotators manually labeled all images using a custom interface, focusing on challenging areas with transparent surfaces, including clean, highly transparent materials, cluttered backgrounds, and strong reflections. They can annotate relative depth by drawing a monotonic depth line, along which depth increases consistently. Sampling points along the line will generate relative depth tuples such as pairs, triplets, or quadruplets. When no clear monotonic structure existed, the annotators used a simpler approach by selecting a reference point instead. In both ways, additional points could be placed in front of (smaller depth) or behind (larger depth) the whole depth line or the reference point, creating tuples of relative depth relationships across different surfaces. Each line and point was also labeled with a layer ID, to specify which layer was being annotated. An example of this process is shown in Fig. 4.
Because multi-layer depth is not a familiar concept in daily life, achieving precise annotations through crowdsourcing is impractical. Therefore, we chose to annotate all images ourselves to ensure data quality. In total, we annotated 1.7M pairs, 4.2M triplets, and 4.2M quadruplets.
@@
Multi-head 69.13 44.85 27.33 76.21 47.55 27.19 65.67 39.94 24.23 65.22 44.26 29.24 62.31 41.45 28.19 69.44 52.56 45.79
Index Concat 69.95 46.42 27.28 78.29 50.07 27.38 66.96 41.85 24.47 65.22 47.39 33.42 69.92 50.92 36.00 82.90 80.15 77.64
Recurrent 70.23 46.37 27.27 77.77 49.08 27.52 68.10 44.47 26.13 65.90 44.26 29.24 62.31 41.45 28.19 70.50 66.56 52.61
@@
ZoeDepth Unidepth V2 GeoWizard Marigold MiDaS MoGe Metric3D V2 DA DA V2 Depth Pro Metric3D V2 ft.
P 74.25 77.03 81.39 82.59 76.61 76.76 80.31 78.02 85.34 87.39 89.53
T 58.56 62.15 66.29 68.35 62.05 63.99 65.43 62.95 74.44 76.29 81.71
Q 52.73 56.85 52.43 55.89 58.54 58.92 55.14 58.88 70.43 69.46 75.20
@@ 7.1 Multi-Layer Depth Baseline Evaluation
There is no clear evidence suggesting the superiority of any particular design. However, one notable observation is that the Layer Index Concatenation method performs exceptionally well on layer 7. This could be because, for larger layer IDs, the model primarily needs to learn how to see through transparent surfaces, and the Index Concat approach provides the most effective prompt for this task.
@@ 7.2 Single-Layer Depth Experiments
$\text{Depth Anything}^{\ddagger}$ 0.73 2.23 0.79 0.89 2.32 5.33 0.39 0.59 1.80 5.30 0.49 0.66 1.75 5.06 0.65 0.75
[…]
@@
Table 3: Representative depth methods evaluated on synthetic validation set. Best scores are in bold. Second best underlined. §$\S$: Metric3D V2 predictions are scaled using ground-truth camera intrinsics. $\dagger$: MoGe is inherently a scale-invariant method, but we estimate an additional global shift for easier comparison with other affine-invariant methods. $\ddagger$: The predictions from MiDas, Depth Anything, and Depth Anything V2 are aligned in disparity space. $\P$: ZoeDepth, UniDepth, Metric3D V2, and Depth Pro are metric depth methods but are evaluated in affine-invariant setting for a fair comparison.
We evaluate ten state-of-the-art depth estimation methods on our real-world benchmark, including Depth Anything (DA) [48], Depth Anything V2 (DA V2) [49], Depth Pro [4], ZoeDepth [3], Unidepth V2 [27], GeoWizard [15], Marigold [19], MiDaS V3.1 [31], MoGe [42], and Metric3D V2 [17]. To assess the effectiveness of our synthetic dataset, we also evaluate a depth model fine-tuned on our synthetic data. We choose Metric3D V2 [17] for fine-tuning, as it provides publicly available fine-tuning code. See Sec. 9.3 for details. Since existing depth models only perform single-layer depth estimation, we evaluate them exclusively on the Layer 1 subset, where models are only required to predict depth for the frontmost layer. Similarly, the Metric3D V2 fine-tuning is conducted solely on the Layer 1 ground truth of our synthetic dataset.
@@
Results are shown in Tab. 3. For metric depth, Metric3D and DepthPro achieve the best results, while for affine-invariant depth, DepthPro performs the best.
Overall, all methods show significantly higher errors on our dataset compared to widely used depth benchmarks, particularly on transparent regions, highlighting the challenges of handling transparency in depth estimation. Notably, the Adapted Layer strategy exhibits errors on transparent objects than the First or Last Layer strategies. This aligns with our observation that existing methods often struggle to disentangle depth information from multiple layers. Rather than accurately predicting each layer, they tend to produce depth estimates that fluctuate between different depth layers at transparent regions. This observation further highlights the importance of multi-layer depth estimation, as the task inherently encourages models to disentangle conflicting multi-layer visual features more effectively.
@@ 8 Conclusion
We propose a new task, multi-layer depth estimation for transparent objects, and introduce LayeredDepth, a real-world depth benchmark and a procedural synthetic data generator designed for the task. We believe our dataset will drive progress in this field.
@@ 9 Appendix
9.1 Benchmark Details
@@
All images and validation annotations in our benchmark are released under the CC0. Test annotations in the benchmark will be withheld for use on a public evaluation server. To validate our approach, we manually annotated 30 synthetic images with known depth ground truth. Our annotations matched the ground truth in 98% of cases, demonstrating the reliability. In total, we annotated 1,500 images, with 300 allocated for validation and 1,200 for testing. Our annotations include 5406 monotonic depth lines and 38392 relative depth points across 7 distinct layers, from which we sampled 1.7M pairs, 4.2M triplets, and 4.2M quadruplets.
9.2 Baseline Training
@@
We train all baseline models from scratch on our synthetic dataset for 100 epochs. During each training step, a random layer is selected as the prediction target. To provide richer supervision, we utilize snapped layered depth: if a pixel lacks ground-truth depth at layer $i$, it inherits the depth value from layer $i-1$. In the Recurrent method, we use the ground-truth depth from the previous layer as input during training, and the model’s output during inference. For optimization, we use the Scale-Invariant Logarithmic loss [13].
9.3 Evaluation and Fine-tuning
@@
All methods are evaluated using a single NVIDIA RTX 3090 GPU. When assessed on the synthetic validation dataset, both the ground-truth values and predictions are clipped to the range $(0.001,30)$. Fine-tuning for Metric3D V2 [17] is performed using their publicly available code, with training for 100,000 steps.

### v2

@@ 1 Introduction
Our real-world benchmark consists of 1,500 images of transparent objects collected from diverse environments, including households, retail spaces, restaurants, laboratories, urban environments, and art installations, under various lighting conditions. Since ground-truth numerical depth cannot be accurately obtained for transparent objects, let alone multi-layer depth, we turn to relative depth annotations instead. Relative depth provides rich information about 3D structures and serves as an effective evaluation metric. Moreover, human annotators excel at determining relative depth, as they can reliably judge which of two points is closer to the camera and provide accurate annotations. In total, we generate 14.2M tuples for relative depth annotations. A gallery of our benchmark along with sample annotations is shown in Fig. 1. All images and validation annotations are available under CC0. Our benchmark is highly challenging for state-of-the-art depth estimation methods [50, 51, 4, 16, 20, 3, 27, 44, 18, 32], even when evaluated on the simplified task of predicting only the first visible layer. For example, Metric3D V2 [18] achieves just 55.14% quadruplet accuracy, while Depth Anything V2 [51], the most accurate among them, reaches only 70.43%.
@@
  • We propose a real-world multi-layer depth benchmark LayeredDepth for transparent objects, consisting of 1,500 CC0 images of diverse scenes and 14.2M relative depth tuples. Our evaluation of state-of-the-art depth estimation methods on our benchmark reveals that they struggle significantly with transparent objects.
@@ 4.2 Annotation Acquisition
The annotators manually labeled all images using a custom interface, focusing on challenging areas with transparent surfaces, including clean, highly transparent materials, cluttered backgrounds, and strong reflections. They can annotate relative depth by drawing a monotonic depth line, along which depth increases consistently. Sampling points along the line will generate relative depth tuples such as pairs, triplets, or quadruplets. When no clear monotonic structure existed, the annotators used a simpler approach by selecting a reference point instead. In both ways, additional points could be placed in front of (smaller depth) or behind (larger depth) the whole depth line or the reference point, creating tuples of relative depth relationships across different surfaces. Each line and point was also labeled with a layer ID, to specify which layer was being annotated. An example of this process is shown in Fig. 4. To penalize models that generate excessive layers and false positives, we also include 10% manually-labeled fake tuples representing nonexistent scenarios (e.g., a tuple labeled as layer 2 in a non-transparent region). During relative depth evaluation, these fake tuples will be marked incorrect if the model predicts any depth at those points.
Because multi-layer depth is not a familiar concept in daily life, achieving precise annotations through crowdsourcing is impractical. Therefore, we chose to annotate all images ourselves to ensure data quality. In total, we annotated 2.5M pairs, 5.9M triplets, and 5.8M quadruplets.
@@
Multi-head 63.42 42.55 25.97 74.72 46.13 26.38 65.66 39.93 24.21 56.58 43.36 30.77 52.17 38.66 29.94 38.40 35.65 33.10
Index Concat 64.46 44.00 26.00 76.70 48.37 26.46 66.95 41.84 24.45 55.85 41.53 30.06 55.36 42.69 31.57 39.66 45.18 45.08
Recurrent 62.36 41.88 24.64 73.51 45.27 25.35 68.08 44.46 26.12 49.47 31.53 21.25 45.51 30.06 21.29 30.09 34.10 27.59
@@
ZoeDepth UniDepth GeoWizard Marigold MiDaS MoGe Metric3D DA V1 Depth4ToM Tosi et al. DA V2 Depth Pro Metric3D ft.
P 74.25 77.03 81.39 82.59 76.61 76.76 80.31 78.02 84.44 84.72 85.34 87.39 89.53
T 58.56 62.15 66.29 68.35 62.05 63.99 65.43 62.95 74.50 73.80 74.44 76.29 81.71
Q 52.73 56.85 52.43 55.89 58.54 58.92 55.14 58.88 70.61 70.10 70.43 69.46 75.20
@@ 7.2 Single-Layer Depth Experiments
$\text{DA}^{\ddagger}$ 0.73 2.23 0.79 0.89 2.32 5.33 0.39 0.59 1.80 5.30 0.49 0.66 1.75 5.06 0.65 0.75
$\text{DA V2}^{\ddagger}$ 0.40 1.51 0.83 0.92 1.24 3.19 0.55 0.74 1.06 3.42 0.45 0.66 0.96 3.01 0.73 0.83
$\text{Depth4ToM}^{\ddagger}$ 0.41 1.08 0.67 0.86 0.70 1.42 0.45 0.69 0.53 1.77 0.41 0.64 0.44 1.29 0.66 0.81
Tosi et al. 0.34 1.03 0.77 0.90 0.65 1.50 0.54 0.74 0.48 1.79 0.50 0.70 0.39 1.32 0.74 0.86
@@
Table 3: Representative depth methods evaluated on synthetic validation set. Best scores are in bold. Second best underlined. $\S$: Metric3D V2 predictions are scaled using ground-truth camera intrinsics. $\dagger$: MoGe is inherently a scale-invariant method, but we estimate an additional global shift for easier comparison with other affine-invariant methods. $\ddagger$: The predictions from MiDas, Depth Anything, Depth Anything V2, and Depth4ToM are aligned in disparity space. $\P$: ZoeDepth, UniDepth, Metric3D V2, and Depth Pro are metric depth methods but are evaluated in affine-invariant setting for a fair comparison.
We evaluate ten state-of-the-art depth estimation methods on our real-world benchmark, including Depth Anything (DA) [50], Depth Anything V2 (DA V2) [51], Depth Pro [4], ZoeDepth [3], Unidepth V2 [28], GeoWizard [16], Marigold [20], MiDaS V3.1 [32], MoGe [44], and Metric3D V2 [18], and two methods tailored for transparent objects, including Tosi et al. [40] and Depth4ToM [12]. To assess the effectiveness of our synthetic dataset, we also evaluate a depth model fine-tuned on our synthetic data. We choose Metric3D V2 [18] for fine-tuning, as it provides publicly available fine-tuning code (see supplementary for details). Since existing depth models only perform single-layer depth estimation, we evaluate them exclusively on the Layer 1 subset, where models are only required to predict depth for the frontmost layer. Similarly, the fine-tuning is conducted solely on the Layer 1 of our synthetic dataset.
@@
Results are shown in Tab. 3. For metric depth, Metric3D and DepthPro achieve the best results, while for affine-invariant depth, DepthPro performs the best. Overall, all methods show significantly higher errors on our dataset compared to widely used depth benchmarks, particularly on transparent regions, highlighting the challenges of handling transparency in depth estimation. Notably, the Adapted Layer strategy exhibits errors on transparent objects than the First or Last Layer strategies. This aligns with our observation that existing methods often struggle to disentangle depth information from multiple layers. Rather than accurately predicting each layer, they tend to produce depth estimates that fluctuate between different depth layers at transparent regions. This observation further highlights the importance of multi-layer depth estimation, as the task inherently encourages models to disentangle conflicting multi-layer visual features more effectively.
@@ 8 Conclusion
We introduce LayeredDepth, a real-world depth benchmark and a procedural synthetic data generator designed for multi-layer depth estimation of transparent objects. We believe our dataset will drive progress in this field.
@@ 9 Appendix
Method All Mixed Layer 1 Layer 3 Layer 5 Layer 7
P T Q P T Q P T Q P T Q P T Q P T Q
Multi-head (NeWCRFs) 63.42 42.55 25.97 74.72 46.13 26.38 65.66 39.93 24.21 56.58 43.36 30.77 52.17 38.66 29.94 38.40 35.65 33.10
Index Concat (NeWCRFs) 64.46 44.00 26.00 76.70 48.37 26.46 66.95 41.84 24.45 55.85 41.53 30.06 55.36 42.69 31.57 39.66 45.18 45.08
Recurrent (NeWCRFs) 62.36 41.88 24.64 73.51 45.27 25.35 68.08 44.46 26.12 49.47 31.53 21.25 45.51 30.06 21.29 30.09 34.10 27.59
Multi-head (DA v2) 80.72 70.67 61.68 89.37 72.98 62.23 85.80 74.34 65.62 73.75 62.68 59.19 66.19 59.29 54.75 57.39 50.78 48.09
Method Layer 1 Layer 3 Layer 5 Layer 7
[…]
Multi-head (NeWCRFs) 17.97 23.41 83.08 93.14 15.90 47.39 80.15 93.89 14.58 47.51 81.48 94.12 16.27 58.78 80.44 93.92
Index Concat (NeWCRFs) 17.26 23.70 83.02 93.27 16.25 48.19 79.63 93.72 15.03 48.11 80.61 93.97 16.49 59.03 80.26 93.76
Recurrent (NeWCRFs) 17.23 25.53 81.89 93.37 17.61 51.79 76.39 92.53 17.25 52.53 76.54 92.20 18.35 61.98 77.04 92.40
Multi-head (DA v2) 8.24 15.21 93.43 97.73 10.68 37.97 88.76 96.83 11.15 40.67 87.51 96.40 13.84 54.63 84.79 95.13
Table 5: Multi-layer baseline methods evaluated on our synthetic validation set. Values are scaled by 100 for clearer comparison. Best scores are in bold. Second best underlined.
9.1 Additional Details
9.1.1 Benchmark
@@
All images in our benchmark are released under the CC0 license. We will make the dataset publicly available; however, a subset of the ground truth annotations for the real-world images will be withheld for use on a public evaluation server. To validate our approach, we manually annotated 30 synthetic images with known depth ground truth. Our annotations matched the ground truth in 98% of cases, demonstrating the reliability. In total, we annotated 1,500 images, with 300 allocated for validation and 1,200 for testing. Our annotations include 5,406 monotonic depth lines and 38,392 relative depth points across 7 distinct layers, plus 3,011 fake depth lines and 6,025 fake relative depth points.
9.1.2 Baseline Training
@@
We train all NeWCRFs baseline models from scratch on our synthetic dataset for 100 epochs. For the Multi-head strategy, we set the number of output depth channels from 1 to 4 to predict multi-layer depth. For the Index Concat strategy, we concatenate the image with an index channel of size $H\times W$, where all values are set to the corresponding layer index, and set the number of input channels to 4 (RGB plus Index). For the Recurrent strategy, we concatenate the image with the previous depth output and also set the number of input channels to 4 (RGB plus Depth). During each training step, a random layer is selected as the prediction target. To provide richer supervision, we utilize snapped layered depth: if a pixel lacks ground-truth depth at layer $i$, it inherits the depth value from layer i1i-1. For optimization, we use the Scale-Invariant Logarithmic loss [14].
9.1.3 Evaluation and Fine-tuning
@@
All methods are evaluated using a single NVIDIA RTX 3090 GPU. When assessed on the synthetic validation dataset, both the ground-truth values and predictions are clipped to the range $(0.001,30)$. Fine-tuning for Metric3D V2 [18] is performed using the publicly available code, with training for 100,000 steps.
9.2 Additional Results
9.2.1 Fine-tuning DepthAnything V2
The proposed multi-layer depth baseline method can be applied to any existing depth backbone. To demonstrate how recent models can benefit from our dataset, we further fine-tuned DepthAnything V2 [51] for multi-layer depth estimation using the multi-head strategy. To achieve this, we replicate the depth head four times, enabling the model to predict multi-layer depth.
Results evaluated on the real world benchmark are shown in Tab. 4. Results evaluated on the synthetic validation set are shown in Tab. 5. Qualitative comparison on real benchmark can be found in Fig. 9 and Fig. 10. Qualitative comparison on synthetic validation set can be found in Fig. 11. These results show that DepthAnything V2 not only aligns more closely with the training domain but also generalizes exceptionally well after fine-tuning solely on our synthetic data. This underscores the power of our synthetic data generator for multi-layer depth estimation and transparent-object understanding. Nevertheless, some regions around object boundaries and in deeper layers still exhibit artifacts, leaving rooms for improved model designs in future work.
9.2.2 Additional Qualitative Results
To help readers better visually assess the effectiveness of our datasets, we provide additional qualitative results on real world benchmark and synthetic validation set, which can be found in Fig. 12 and Fig. 13, respectively.
Figure 9: Additional qualitative comparison of multi-layer depth baselines on real benchmark.
Figure 10: Additional qualitative comparison of multi-layer depth baselines on real benchmark.
Figure 11: Additional qualitative comparison of multi-layer depth baselines on synthetic validation set.
Figure 12: Additional qualitative comparison of state-of-the-art single layer depth methods on our benchmark.
Figure 13: Additional qualitative comparison of state-of-the-art single layer depth methods on our synthetic validation set.